OpenStudy (anonymous):

What is the x-intercept of the equation 6x+8y+9=0

OpenStudy (agreene):

You can solve with respect to y and it will place this in y=mx+b form, giving you the x-intercept. 6x+8y+9=0 8y=-9-6x y=-3/4x-9/8 therefore your x-intercept will be -9/8 :)

OpenStudy (anonymous):

Thank you :

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20081029234449AAWwhLy When y is zero, 6x = -9, x = -3/2 A plot of 6x+8y+9=0 is attached.